# An Efficient Global Optimization Algorithm with Adaptive Estimates of the Local Lipschitz Constants

### v3

@@ Abstract
In this work, we present a new deterministic partition-based Global Optimization (GO) algorithm that uses estimates of the local Lipschitz constants associated with different sub-regions of the domain of the objective function. The estimates of the local Lipschitz constants associated with each partition are the result of adaptively balancing the global and local information obtained so far from the algorithm, given in terms of absolute slopes. We motivate a coupling strategy with local optimization algorithms (both gradient-based and derivative-free) to accelerate the convergence speed of the proposed approach. In the end, we compare our approach HALO (Hybrid Adaptive Lipschitzian Optimization) with respect to popular GO algorithms using hundreds of test functions. From the numerical results, the performance of HALO is very promising and can extend our arsenal of efficient procedures for attacking challenging real-world GO problems. The Python code of HALO is publicly available on GitHub. [^1] https://github.com/dannyzx/HALO
@@ 1 Introduction
Global optimization (GO) is a critical field in operations research, focusing on methodologies and algorithms for finding the global optimal solution to Problem 1. The pursuit of finding the global optimum covers many scientific disciplines due to its fundamental importance. Extensive academic research has been dedicated to GO, resulting in a plethora of proposed approaches. Stochastic and evolutionary algorithms, such as controlled random search [1], genetic algorithm [2], simulated annealing [3], particle swarm optimization [4], and covariance matrix adaptation evolutionary strategy (CMA-ES) [5], have proven effective in numerous real-world applications. Another class of GO algorithms involves constructing response surfaces of the objective function using statistical models like gaussian processes [6, 7] or radial basis functions [8, 9]. These methodologies exhibit significant performance improvements in scenarios where evaluating the objective function $f$ is computationally expensive.
Lipschitz optimization, extensively studied and developed over the past decades [10, 11, 12, 13, 14], constitutes another class of GO algorithms. In Lipschitz optimization, the objective function $f$ is assumed to be Lipschitz continuous on the feasible domain $\mathcal{D}$. This implies the existence of a constant $0<L<\infty$ such that
$|f(\mathbf{x})-f(\bar{\mathbf{x}})|\leq L||\mathbf{x}-\bar{\mathbf{x}}||\qquad\forall\mathbf{x},\bar{\mathbf{x}}\in\mathcal{D}$ (2)
where $L$ represents the global Lipschitz constant of the function $f$, and |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm (alternative norms can be utilized, as explored in [15]). The requirement for $f$ to be Lipschitz-continuous is not overly restrictive, as this class encompasses a wide range of functions. For instance, any continuously differentiable function defined within a convex and compact set is Lipschitz continuous [14], with $L$ corresponding to the maximum norm of the gradient $L=\max_{\mathbf{x}\in\mathcal{D}}||\nabla f(\mathbf{x})||$. Leveraging the Lipschitz continuity of the objective function $f$ is particularly useful, as it allows for the computation of valid lower bounds across the domain $\mathcal{D}$. By knowing the function value at a point $f(\bar{\mathbf{x}})$, it becomes possible to compute the lower bound for any $\mathbf{x}\in\mathcal{D}$ that satisfies the following inequality
$f(\bar{\mathbf{x}})-L||\mathbf{x}-\bar{\mathbf{x}}||\leq f(\mathbf{x})$ (3)
@@ 2 Related Work
During the years, researchers developed Lipschitz optimization methods as in [16, 17] that given the value of the Lipschitz constant $L$, can produce valid lower bounds of the objective function $f$ and driving the numerical procedure towards the global minimizers.
@@
The difficulty is when the value $L$ of the Lipschitz constant is not available in advance such as in the majority of real-world applications. To effectively overcome this issue, an estimation of it is usually carried out given the information acquired by the optimization procedure during its iterations as in [12] and using a diagonal partition scheme in [18, 19, 20].
@@
One popular method that circumvents the need for estimating the Lipschitz constant $L$ is the DIRECT (DIviding RECTangles) algorithm [21]. This derivative-free, partition-based algorithm operates under the determination of potentially optimal rectangles, where any value of the global Lipschitz constant $L$ within the range $(0,\infty)$ is implicitly considered during the selection of the partitions to explore. DIRECT possesses several noteworthy features, including a simple partition scheme, an everywhere-dense property that ensures convergence towards the global optimal solution as the algorithm iterates indefinitely, and a minimal number of hyperparameters. In fact, DIRECT has only one hyperparameter and has exhibited robustness in practical applications [21]. These characteristics have inspired the development of several similar algorithms based on the concept of potentially optimal partitions [22, 23, 24, 25, 26, 27, 28, 29].
@@
Other noteworthy algorithms in the Lipschitz optimization class focus on estimating the local Lipschitz constant $L_{i}$ for each sub-region $\mathcal{D}_{i}$ of the domain $\mathcal{D}$. In [30], an algorithm is proposed for optimizing univariate functions, while for multivariate functions, a dimensionality reduction method employing Peano curves is utilized [31, 32]. The estimation of local Lipschitz constants is employed in a diagonal partition-based algorithm described in [33]. In [34], the computation of local Lipschitz constants is integrated into a scheme that reduces the initial multidimensional problem to a set of recursively and adaptively connected univariate subproblems, as further explored in [35]. A similar concept is proposed when the first derivative is available in [36]. However, it is worth noting that the performance of these approaches is highly susceptible to a reliability parameter that users must define beforehand, as emphasized in [33, 34, 37, 38]. Similar approaches have been developed specifically for univariate objective functions [39, 40, 41, 42].
@@
Furthermore, successful hybridizations have been achieved by combining some of the aforementioned approaches with local optimization methods. For instance, the DIRECT algorithm has been hybridized with a truncated Newton method in [43] and a coordinate descent approach in [44], resulting in improved performance, as also observed in the methodology presented in [24].
@@ 3 Main Contribution
We propose a criterion for selecting the most promising partitions based on the information provided by the local Lipschitz constants. This criterion deviates from the concept of potentially optimal hyperrectangles employed in the DIRECT algorithm. Furthermore, we establish the asymptotic convergence of our algorithm called HALO (Hybrid Adaptive Lipschitzian Optimization) to a global minimum, ensuring the everywhere dense property. To expedite the convergence speed of HALO towards stationary points, we introduce a simple coupling strategy with two different local optimization algorithms: a gradient-based and a derivative-free optimization approach presented respectively in [45] and [46].
@@
We evaluate the performance of HALO by comparing it to popular GO algorithms such as DIRECT [21], CMA-ES [5], L-SHADE [47], and the hybridization of DIRECT known as DIRMIN [44]. To assess the algorithms, we conduct an extensive numerical campaign using well-known test functions collected from [48], as well as two additional test function generators specifically designed for GO problems presented in [49, 50]. In total, we consider over one thousand test functions.
@@ 20       $\mathcal{D}_{i_{k}^{\star}}\leftarrow\bigcup\limits_{j=1}^{J}\mathcal{D}_{i_{k}^{\star}}^{j}$
The Main function, which takes the maximum number of iterations (MaxIter) and the box domain ($\mathcal{D}$) as input, serves as the main entry point of the algorithm. In line 3, the algorithm begins with a loop over the iterations $k$. Within this loop, the Selection function is called in line 4 to select partitions based on certain criteria which uses the information about the Lipschitz constant of each partition (not yet specified). The selected partitions are represented by the subset of indices $\mathcal{I}^{\star}_{k}$ obtained from the set of indices $\mathcal{I}_{k}$, where each index $i_{k}$ corresponds to a partition $\mathcal{D}_{i_{k}}$ at iteration $k$. Lines 5 and 6 contain two nested ’for’ loops. The outer loop iterates over each index $i^{\star}_{k}$ of the selected partitions, while the inner loop (line 6) performs $J$ function evaluations within each partition. The function evaluations are carried out at points $\mathbf{x}^{j}_{i^{\star}_{k}}$ sampled from the partition $\mathcal{D}_{i^{\star}_{k}}$. The corresponding function values $f(\mathbf{x}^{j}_{i^{\star}_{k}})$ are computed using the Fun_Eval. Once the inner loop completes, the set $\mathcal{D}_{i^{\star}_{k}}$ is divided into $J$ non-overlapping subsets (line 16), denoted as $\mathcal{D}_{i_{k}^{\star}}^{j}$ for $j=1,\dots,J$. Furthermore, the local Lipschitz constant around every partition is estimated at line 10 and used in the Selection function. In line 17, the set of indices $\mathcal{I}_{k}$ is updated to include the new indices resulting from the partitioning step.
@@ 4.1.1 Adaptive Estimation of the Local Lipschitz Constants
In Lipschitz optimization, it is common practice to obtain information about the Lipschitz constant $L$ and compute lower bounds of the objective function $f$ to guide the algorithm towards regions where these bounds are lower. However, relying solely on the global Lipschitz constant estimates may not accurately reflect the behavior of the objective function in specific regions of the domain $\mathcal{D}$.
Consider a scenario where the objective function $f$ is mostly flat, except for a small region within $\mathcal{D}$ where it exhibits highly chaotic behavior (i.e., a higher Lipschitz constant). In such cases, the lower bounds computed around the gentler regions of $\mathcal{D}$ would poorly estimate the true lower bounds because the global Lipschitz constant represents the overall maximum variation of $f$. Therefore, it is more informative to estimate the local Lipschitz constant $\tilde{L}_{i}$ associated with each partition $\mathcal{D}_{i}$ instead of relying entirely on the global Lipschitz constant $L$.
@@
Given $\mathcal{D}_{i_{k}}$ as a partition of the feasible space at iteration $k$, and let
$\alpha_{i_{k}}=\frac{||\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||}{\sqrt{N}},\quad\alpha_{i_{k}}\in(0,1),\quad\tilde{L}_{k}=\max_{i_{k}\in\mathcal{I}_{k}}||\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}||$ (4)
the estimation of the local Lipschitz constant $L_{i_{k}}$ is given by
$\tilde{L}_{i_{k}}=\alpha_{i_{k}}\tilde{L}_{k}+\left(1-\alpha_{i_{k}}\right)||\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}||$ (5)
In Eq. 5, the term $\alpha_{i_{k}}$ is the ratio between the diagonal of the partition $\mathcal{D}_{i_{k}}$ and the main diagonal of the feasible set $\mathcal{D}$ (normalized to a unit hypercube). As the diagonal of the partition $\mathcal{D}_{i_{k}}$, denoted by $||\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||$, decreases, the local information in the neighborhood of $\mathcal{D}_{i_{k}}$ becomes more precise, and the estimate of the local Lipschitz constant is primarily influenced by the norm of the gradient approximation at $\mathbf{x}_{i_{k}}$, namely ~$||\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}||$. On the other hand, if the diagonal of the partition $\mathcal{D}_{i_{k}}$ is approximately equal to the maximum diagonal, the local information may not be reliable, and the estimate of the local Lipschitz constant relies more on the estimate of the global Lipschitz constant $\tilde{L}_{k}$, which represents the maximum absolute variation observed during the iterations of the algorithm.
@@ 1. Criterion  1
    if for $i_{k}\in\mathcal{I}_{k}$, then
    $f(\mathbf{x}_{i_{k}})-\tilde{L}_{i_{k}}\frac{||\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||}{2}\leq f(\mathbf{x}_{j_{k}})-\tilde{L}_{j_{k}}\frac{||\mathbf{u}_{j_{k}}-\mathbf{l}_{j_{k}}||}{2}\qquad\forall j\in\mathcal{I}_{k}\$ (6)
@@ 3. Criterion  3
    if for $i_{k}\in\mathcal{I}_{k}^{\text{max}}$ and
    $f(\mathbf{x}_{i_{k}})-\tilde{L}_{i_{k}}\frac{||\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||}{2}\leq f(\mathbf{x}_{j_{k}})-\tilde{L}_{j_{k}}\frac{||\mathbf{u}_{j_{k}}-\mathbf{l}_{j_{k}}||}{2}\qquad\forall j\in\mathcal{I}_{k}^{\text{max}}$ (8)
    where $\mathcal{I}_{k}^{\text{max}}=\{i_{k}\in\mathcal{I}_{k}:\frac{||\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||}{2}=\max_{i_{k}\in\mathcal{I}_{k}}\frac{||\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||}{2}\}$.
@@
  • Criterion  1: This criterion selects the partition $\mathcal{D}_{i_{k}}$ that achieves the lowest lower bound, considering the information collected from the local Lipschitz constants. The lower bound is computed as $f(\mathbf{x}_{i_{k}})-\tilde{L}_{i_{k}}||\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||/2$, where $f(\mathbf{x}_{i_{k}})$ is the function value at the centroid of $\mathcal{D}_{i_{k}}$, $\tilde{L}_{i_{k}}$ is the estimated local Lipschitz constant, and $||\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||/2$ is the distance from the center to the vertices of the partition. This criterion aims to select partitions that provide the most promising lower bounds based on the local behavior of the objective function.
@@
In Fig. 1, we can compare the selection criteria of HALO with those of the DIRECT algorithm, specifically targeting the potentially optimal partitions. We used the synthetic function generator presented in [49] for this particular example. Notably, HALO follows distinct trajectories from DIRECT, as none of its selected partitions align with the potentially optimal rectangles. In fact, HALO has the ability to choose rectangles that can be arbitrarily distant from the potentially optimal partitions of DIRECT. This allows HALO to proactively explore interesting partitions by leveraging its knowledge of the local Lipschitz constants. As illustrated in Fig. 1, HALO efficiently identifies the partition (indicated by the white star) where the global minimum is located since is the one that satisfies Criterion  1, while DIRECT experiences a delay in identifying the optimal partition.

### v4

@@ Abstract
In this work, we present a new deterministic partition-based global optimization algorithm, HALO (Hybrid Adaptive Lipschitzian Optimization), which uses estimates of the local Lipschitz constants associated with different sub-regions of the objective function’s domain to compute lower bounds and guide the search toward global minimizers. These estimates are obtained by adaptively balancing the global and local information collected from the algorithm, based on absolute slopes. HALO is hyperparameter-free, eliminating the need for manual tuning, and it highlights the most important variables to help interpret the optimization problem. We also introduce a coupling strategy with local optimization algorithms, both gradient-based and derivative-free, to accelerate convergence. We compare HALO with popular global optimization algorithms on hundreds of test functions. The numerical results are very promising and demonstrate that HALO can expand our arsenal of efficient procedures of efficient procedures for challenging real-world black-box optimization problems. The Python code of HALO is publicly available on GitHub. [^1] https://github.com/dannyzx/HALO
@@ 1 Introduction
Global optimization (GO) is a critical field in operations research, focusing on methodologies and algorithms for finding the global optimal solution to Problem 1. The pursuit of finding the global optimum covers many scientific disciplines due to its fundamental importance. Extensive academic research has been dedicated to GO, resulting in a wide range of proposed approaches. Stochastic and evolutionary algorithms, such as controlled random search [price1977controlled], genetic algorithm [davis1991handbook], simulated annealing [johnson1989optimization], particle swarm optimization [kennedy1995particle], and covariance matrix adaptation evolutionary strategy (CMA-ES) [hansen2001completely], have proven effective in numerous real-world applications. Another class of GO algorithms involves constructing response surfaces of the objective function using statistical models like gaussian processes [jones1998efficient, mockus1978application] or radial basis functions [gutmann2001radial, regis2007stochastic]. These methodologies exhibit good performance in scenarios where evaluating the objective function $f$ is computationally expensive and evaluation budget is limited.
Lipschitz optimization, extensively studied and developed over the past decades [horst2013global, horst2013handbook, sergeyev2006global, torn1989global, pinter1995global], constitutes another class of GO algorithms. In Lipschitz optimization, the objective function $f$ is assumed to be Lipschitz continuous on the feasible domain $\mathcal{D}$. This implies the existence of a constant $0<L<\infty$ such that
[…]
@@ 2 Related Work
Over the years, researchers have developed various Lipschitz optimization methods (e.g., [shubert1972sequential, mladineo1986algorithm]) which, given the value of the Lipschitz constant $L$, can produce valid lower bounds for the objective function $f$, guiding the numerical procedure toward the global minimizers.
@@
The difficulty is when the value $L$ of the Lipschitz constant is not available in advance such as in the majority of real-world applications. To effectively overcome this issue, an estimation of it is usually carried out given the information acquired by the optimization procedure during its iterations as in [sergeyev2006global] and using a diagonal partition scheme in [kvasov2012lipschitz, sergeyev2017deterministic, sergeyev2015deterministic2].
@@
One popular method that circumvents the need for estimating the Lipschitz constant $L$ is the DIRECT (DIviding RECTangles) algorithm [jones1993-JOTA]. This derivative-free, partition-based algorithm operates under the determination of potentially optimal rectangles, where any value of the global Lipschitz constant $L$ within the range $(0,\infty)$ is implicitly considered during the selection of the partitions to explore. DIRECT possesses several noteworthy features, including a simple partition scheme, an everywhere-dense property that ensures convergence towards the global optimal solution as the algorithm iterates indefinitely, and a minimal number of hyperparameters. In fact, DIRECT has only one hyperparameter and has exhibited robustness in practical applications [jones1993-JOTA]. These characteristics have inspired the development of several similar algorithms based on the concept of potentially optimal partitions [gablonsky2001, mockus_2011, paulavivcius2020globally, mockus2017application, liu2015mrdirect, liu2017improving, lera2015deterministic, paulavivcius2014simplicial].
@@
Other noteworthy algorithms in the Lipschitz optimization class focus on estimating the local Lipschitz constant $L_{i}$ for each sub-region $\mathcal{D}_{i}$ of the domain $\mathcal{D}$. In [sergeyev1995information], an algorithm is proposed for optimizing univariate functions, while for multivariate functions, a dimensionality reduction method employing Peano curves is utilized [strongin1992algorithms, sergeyev2013introduction]. The estimation of local Lipschitz constants is employed in a diagonal partition-based algorithm described in [kvasov2003local]. In [gergel2015local], the computation of local Lipschitz constants is integrated into a scheme that reduces the initial multidimensional problem to a set of recursively and adaptively connected univariate subproblems, as further explored in [gergel2016adaptive]. A similar concept is proposed when the first derivative is available in [gergel1997global]. However, it is worth noting that the performance of these approaches is highly susceptible to a reliability parameter that users must define beforehand, as emphasized in [kvasov2003local, gergel2015local, jones2020direct, sergeyev2015deterministic]. At the same time, even with a misspecified reliability parameter, these methods can exhibit stronger theoretical convergence guarantees than DIRECT-type methods [Sergeyev01011998, sergeyev2017deterministic]. Similar approaches have been developed specifically for univariate objective functions [sergeyev1998global, sergeyev2021novel, posypkin2022efficient, kvasov2009univariate].
@@
Furthermore, successful hybridizations have been achieved by combining some of the aforementioned approaches with local optimization methods. For instance, the DIRECT algorithm has been hybridized with a truncated Newton method in [liuzzi2010direct] and a coordinate descent approach in [liuzzi2016exploiting], resulting in improved performance, as also observed in the methodology presented in [paulavivcius2020globally].
@@ 3 Main Contribution
We propose a criterion for selecting the most promising partitions based on the information provided by the local Lipschitz constants. This criterion deviates from the concept of potentially optimal hyperrectangles employed in the DIRECT algorithm. Furthermore, we establish the asymptotic convergence of our algorithm called HALO (Hybrid Adaptive Lipschitzian Optimization) to a global minimum, ensuring the everywhere dense property. To expedite the convergence speed of HALO towards stationary points, we introduce a simple coupling strategy with two different local optimization algorithms: a gradient-based and a derivative-free optimization approach presented respectively in [byrd1995limited] and [lucidi2002-COA].
@@
We evaluate the performance of HALO by comparing it to popular GO algorithms such as DIRECT [jones1993-JOTA], CMA-ES [hansen2001completely], L-SHADE [tanabe2014improving], PLOR [mockus2017application] and the hybridization of DIRECT known as DIRMIN [liuzzi2016exploiting]. To assess the algorithms, we conduct an extensive numerical campaign using well-known test functions collected from [Jamil_2013], as well as two additional test function generators specifically designed for GO problems presented in [schoen1993wide, gaviano2003algorithm]. In total, we consider over one thousand test functions.
@@ 20 $\mathcal{D}_{i_{k}^{\star}}\leftarrow\bigcup\limits_{j=1}^{J}\mathcal{D}_{i_{k}^{\star}}^{j}$
The Main function, which takes the maximum number of iterations (MaxIter) and the box domain ($\mathcal{D}$) as input, serves as the main entry point of the algorithm. In line 3, the algorithm begins with a loop over the iterations $k$. Within this loop, the Selection function is called at line 4 to select partitions based on certain criteria, which make use of the Lipschitz constant information of each partition (not yet specified).
The selected partitions are represented by the subset of indices $\mathcal{I}^{\star}_{k}$ obtained from the set of indices $\mathcal{I}_{k}$, where each index $i_{k}$ corresponds to a partition $\mathcal{D}_{i_{k}}$ at iteration $k$. Lines 5 and 6 contain two nested ’for’ loops. The outer loop iterates over each index $i^{\star}_{k}$ of the selected partitions, while the inner loop (line 6) performs $J$ function evaluations within each partition. The function evaluations are carried out at points $\mathbf{x}^{j}_{i^{\star}_{k}}$ sampled from the partition $\mathcal{D}_{i^{\star}_{k}}$. The corresponding function values $f(\mathbf{x}^{j}_{i^{\star}_{k}})$ are computed using the Fun_Eval. Once the inner loop completes, the set $\mathcal{D}_{i^{\star}_{k}}$ is divided into $J$ non-overlapping subsets (line 16), denoted as $\mathcal{D}_{i_{k}^{\star}}^{j}$ for $j=1,\dots,J$. Furthermore, the local Lipschitz constant around every partition is estimated at line 10 and used in the Selection function. In line 17, the set of indices $\mathcal{I}_{k}$ is updated to include the new indices resulting from the partitioning step.
@@ 4.1.1 Adaptive Estimation of the Local Lipschitz Constants
In Lipschitz optimization, it is common practice to obtain information about the Lipschitz constant $L$ and compute lower bounds of the objective function $f$ to guide the algorithm towards regions where these bounds are lower. However, relying solely on the global Lipschitz constant estimates may not accurately reflect the behavior of the objective function in specific regions of the domain $\mathcal{D}$. Beyond global estimates, it is well established that employing local Lipschitz constants $\tilde{L}_{i}$ for each subregion $\mathcal{D}_{i}$ can improve efficiency; see Section 2, for example, the local tuning approaches developed in [sergeyev1998global, sergeyev2017deterministic, sergeyev1995information, strongin2013global, gergel2015local].
To better understand why local estimates can be more informative, consider a scenario where the objective function $f$ is mostly flat, except for a small region within $\mathcal{D}$ where it exhibits highly chaotic behavior (i.e., a much higher Lipschitz constant). In such cases, using a global Lipschitz constant to compute lower bounds across all of $\mathcal{D}$ would lead to poor estimates in the flatter regions. This is because the global constant is dominated by the steep variations in that small chaotic area, overestimating the possible change of $f$ elsewhere and thus weakening the bounds. Therefore, it is more informative to estimate the local Lipschitz constant $\tilde{L}_{i}$ associated with each partition $\mathcal{D}_{i}$ instead of relying entirely on the global Lipschitz constant $L$.
@@
Given a subregion $\mathcal{D}_{i_{k}}$ of the unit hypercube domain $\mathcal{D}$ at iteration $k$, let
[…]
and define the global Lipschitz constant estimate as
[…]
Then we define
[…]
In Eq. 4, the term $\alpha_{i_{k}}$ is the ratio between the diagonal of the partition $\mathcal{D}_{i_{k}}$ and the main diagonal of the feasible set $\mathcal{D}$ (normalized to a unit hypercube). As the diagonal of the partition $\mathcal{D}_{i_{k}}$, denoted by $\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||$, decreases, the local information in the neighborhood of $\mathcal{D}_{i_{k}}$ becomes more precise, and the estimate of the local Lipschitz constant is primarily influenced by the norm of the gradient approximation at $\mathbf{x}_{i_{k}}$, namely $\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}||$. On the other hand, if the diagonal of the partition $\mathcal{D}_{i_{k}}$ is approximately equal to the maximum diagonal, the local information may not be reliable, and the estimate of the local Lipschitz constant relies more on the estimate of the global Lipschitz constant $\tilde{L}_{k}$, which represents the maximum absolute variation observed during the iterations of the algorithm. In this way, our formula (Eq. 4) naturally balances global and local information in a self-adaptive manner, eliminating the need for user-defined parameters as traditional local tuning techniques [sergeyev1998global, sergeyev2017deterministic, sergeyev1995information, strongin2013global, gergel2015local].
@@ 1. Criterion  1
    Given $i_{k}\in\mathcal{I}_{k}$, then
[…]
@@ 3. Criterion  3
    Given $i_{k}\in\mathcal{I}_{k}^{\text{max}}$ and
[…]
@@
[…]
@@
In Fig. 1, we can compare the selection criteria of HALO with those of the DIRECT algorithm, specifically targeting the potentially optimal partitions. We used the synthetic function generator presented in [schoen1993wide] for this particular example. Notably, HALO follows distinct trajectories from DIRECT, as none of its selected partitions align with the potentially optimal rectangles. In fact, HALO has the ability to choose rectangles that can be arbitrarily distant from the potentially optimal partitions of DIRECT. This allows HALO to proactively explore interesting partitions by leveraging its knowledge of the local Lipschitz constants. As illustrated in Fig. 1, HALO efficiently identifies the partition (indicated by the white star) where the global minimum is located since is the one that satisfies Criterion  1, while DIRECT experiences a delay in identifying the optimal partition.